Evaluate the indefinite integral in each item

  1. ∫e2x+1dx
  2. ∫1/(1+ex) dx
  3. ∫e2x/(ex+3) dx
  4. ∫4(102x^2) dx
  5. ∫x2/(x3+5) dx
1
Expert's answer
2021-04-15T06:49:26-0400
  1. "\int e^{2x+1}dx=\frac12\int e^{2x+1}d(2x+1)=\frac12e^{2x+1}+C,C\in{\mathbb{R}}" .
  2. "\int\frac{dx}{e^x+1}". We make the change: "z=e^x". We point out that "z>0". We receive: "\int\frac{dz}{z(z+1)}=\int(\frac{1}{z}-\frac{1}{z+1})dz=ln\,z-ln(z+1)+C=x-ln(e^x+1)+C,C\in{\mathbb{R}}"
  3. "\int\frac{e^{2x}}{e^x+1}dx" . We make the change "z=e^x" and receive: "\int\frac{zdz}{z+1}=\int(1-\frac{1}{z+1})dz=z-ln(z+1)+C=e^x-ln(e^x+1)+C,C\in{\mathbb{R}}"
  4. "\int 4(10^{2x^2})dx=\int4e^{2x^2ln2}dx." The good expression for this integral in terms of elementary functions is not known.
  5. "\int\frac{x^2}{x^3+5}dx=\int\frac{1}{3(x^3+5)}d(x^3+5)=3ln(x^3+5)+C,C\in{\mathbb{R}}."